OpenStudy (anonymous):

Points: 3Question 1 Maggie and Tanya went to a fruit store. Maggie bought 10 oranges and 5 apples for $4. Tanya bought 9 oranges and 3 apples for $3. What is the cost of an orange and an apple? Cost of an orange = $0.10, cost of an apple = $0.40. Cost of an orange = $0.20, cost of an apple = $0.40. Cost of an orange = $0.24, cost of an apple = $0.50. Cost of an orange = $0.10, cost of an apple = $0.50.

OpenStudy (sh3lsh):

Try this. 10x + 5y = 4 9x + 3y = 3 Solve for x and y.

OpenStudy (anonymous):

hm i did

OpenStudy (anonymous):

i did all the steps but when i tried to get the second variable it didnt add up

OpenStudy (anonymous):

i got y= 0.4 where y is apples

OpenStudy (anonymous):

then i put in y and got x = 5

OpenStudy (anonymous):

but the answers arent on the page so its wrong

OpenStudy (anonymous):

cant tell what i got wrong

OpenStudy (sh3lsh):

Let's try substitution. \[3y = 3-9x \] \[y = 1-3x\] \[10x + 5(1-3x) = 4\] \[10x + 5 -15x = 4\] \[-5x = -1\] \[x = 1/5, x= .20\]

OpenStudy (anonymous):

its not up there

OpenStudy (sh3lsh):

\[10(1/5) + 5y = 4, 2 + 5y = 4, 5y = 2, y = 2/5, y = .4 \]

OpenStudy (sh3lsh):

B

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

u thought u meant x=1/5,x=.20 was the answer i see

OpenStudy (anonymous):

i thought

OpenStudy (sh3lsh):

Yeah, sorry. I'm not sure how to make new lines using the equation editor.

OpenStudy (sh3lsh):

Ah! It's two of these \\